Suppose Firm A has a supply curve of and Firm B has a supply curve of How much is the total supply at a price of $4? Tot al supply is thousand units (enter your response as a real number rounded to one decimal place) How much is total supply at a price of $13? Total supply is thousand unts (enter your response as a real number rounded to one deoimal place)
Answer #1

We know market supply is the horizontal summation of individual supply curve

Qa=-2+4=2 and Qb=0.5P=0.5(4)=2

Thus total supply is Qa+Qb=2+2=4thousand

At price=13, Qa=11 and Qb=6.5

thus total supply is 11+6.5=17.5thousand units